# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ 1.2 Main results
1:Input: State $\omega^{\otimes n}$, $s\in\mathbb{N}$, $\eta\in\mathbb{R}$, bases $\mathcal{B}_{\mathcal{A}}=\{Z_{i}\}_{i=1}^{d_{\mathcal{A}}},\mathcal{B}_{\mathcal{A}_{L}}=\{Y_{i}\}_{i=1}^{d^{s}_{\mathcal{A}}},\mathcal{B}_{\mathcal{A}_{R}}=\{X_{i}\}_{i=1}^{d^{s}_{\mathcal{A}}}$
2:Use Tomography$(\omega^{\otimes n},s)$ and produce empirical estimates:
  • $\widehat{\Omega(\mathds{1})}[Y_{j}]=\hat{\omega}_{s}(Y_{j})$
  • $\widehat{\tau\Omega}[X_{i}]=\hat{\omega}_{s}(X_{i})$
  • $\hat{\Omega}(X_{i})[Y_{j}]=\hat{\omega}_{2s}(Y_{j}\otimes X_{i})$
  • $\hat{\Omega}_{Z_{k}}(X_{i})[Y_{j}]=\hat{\omega}_{2s+1}(Y_{j}\otimes Z_{k}\otimes X_{i})$
@@
3:Compute the SVD of $\hat{\Omega}=\hat{U}\hat{D}\hat{W}^{\intercal}$ and truncate the columns of $\hat{U}$ such that we keep the columns corresponding the singular values $\sigma_{i}(\hat{\Omega})\geq\eta$.
4:Compute the estimated realization operators:
@@
5:Reconstruct $\hat{\omega}_{t}$ using
$\hat{\omega}_{t}(Z_{1}\otimes\cdots\otimes Z_{t}):=\hat{\rho}\hat{\mathbb{K}}_{Z_{1}}\cdots\hat{\mathbb{K}}_{Z_{t}}\hat{e}$
for all $Z_{i}\in\mathcal{B}_{\mathcal{A}}$.

### v3

@@ 2.2 Learning algorithm for finitely correlated states
1:Input: State $\omega_{2s+1}^{\otimes n}$, $s\in\mathbb{N}$, $\eta\in\mathbb{R}$, bases $\mathcal{B}_{\mathcal{A}}=\{Z_{i}\}_{i=1}^{d^{2}_{\mathcal{A}}},\mathcal{B}_{L}=\{Y_{i}\}_{i=1}^{d^{2s}_{\mathcal{A}}},\mathcal{B}_{R}=\{X_{i}\}_{i=1}^{d^{2s}_{\mathcal{A}}}$
2:Use Tomography$(\omega_{2s+1}^{\otimes n},s)$ and produce empirical estimates:
  • $\widehat{\Omega(\mathds{1})}[Y_{j}]=\operatorname{Tr}[\hat{\omega}_{s}Y_{j}]$
  • $\widehat{\tau\Omega}[X_{i}]=\operatorname{Tr}[\hat{\omega}_{s}X_{i}]$
  • $\hat{\Omega}(X_{i})[Y_{j}]=\operatorname{Tr}[\hat{\omega}_{2s}(Y_{j}\otimes X_{i})]$
  • $\hat{\Omega}_{Z_{k}}(X_{i})[Y_{j}]=\operatorname{Tr}[\hat{\omega}_{2s+1}(Y_{j}\otimes Z_{k}\otimes X_{i})]$
@@
3:Compute the SVD of $\hat{\Omega}=\hat{U}\hat{D}\hat{O}^{\intercal}$ and truncate the columns of $\hat{U}$ such that we keep the columns corresponding the singular values $\sigma_{i}(\hat{\Omega})\geq\eta/2$.
4:Output the estimated realization operators: